OpenStudy (anonymous):

How to factor 2x^2-5+3

OpenStudy (anonymous):

is it ????\[2x ^{2}-5 + 3\]

OpenStudy (anonymous):

If yes, then it is equal to 2x^2 - 2 = 2 (x^2 - 1) = 2 (x^2 - 1^2) = 2 (x + 1)(x - 1)